Assume that you want to transfer HDTV video of 1000x2000x24 pixels at 30 frames/sec over a network. Available bandwidth of the network is 200 Mbps. How much you must compress the video in order to be able to send it in real time? Calculate the compression ratio.
Answer #1

Number of frames per second = 30

Each frame contains = (1000*2000*24) bits = 48*106 bits = 48 Megabits of data

So, the required bandwidth to transmit the video is = 48*30 Mbps = 1440 Mbps

Since the available bandwidth is 200 Mbps so we can not send this file in real time without compression.

The required minimum compression ratio will be 200:1440 or 1 : 7.2

We can also express the minimum compression ratio as ((200/1440)*100)% = 13.89%

Hope this helps.